# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 6.2 Cyclic group case
Since $\hat{\mathbb{A}}$ is semi-simple, we have $\hat{U}\cong\hat{U}_{1}\oplus\hat{U}_{2}$ as $\hat{\mathbb{A}}-$modules. Then by [16, Theorem 4.16.4], there is an $\mathcal{O}$-pure submodule $U_{1}$ of $U$ such that $\hat{U}_{1}$ is isomorphic to $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$. By [16, Proposition 4.2.6], $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ can be viewed as a submodule of $k\mathop{\otimes}\limits_{\mathcal{O}}U$ which is a projective indecomposable $\bar{\mathbb{A}}$-module. Since $\mathbb{A}$ is a symmetric $\mathcal{O}$-algebra, the socle of $k\mathop{\otimes}\limits_{\mathcal{O}}U$ is simple. So is the socle of $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$. In particular, $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ is indecomposable. By Lemma 6.30, $\hat{U}_{1}]$ is not contained in $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$. ∎
@@
Due to the structure of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ described in Proposition 6.24 (iii), by Lemmas 6.30 and 6.31, the proof is similar to the one of [17, Theorem 11.10.5]. ∎
6.33.
@@
By Theorem 6.20 and [18, Theorem 2.6], the isometry bijection $\Phi_{\mathcal{K}}$ in the commutative diagram (6.23.5) preserves the actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}(L)$. Then the statements on stabilizers of simple $\hat{\mathbb{A}}-$modules are clear. For any $\chi\in\mathcal{M}_{R}$, since $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is coprime to $p$, the same argument in the proof of Proposition 6.15 shows that $\hat{U}_{\chi}$ can be extended to $\hat{\mathbb{A}}_{R_{\chi}}$. For any $\lambda\in\mathrm{Irr}(E)$, since $\hat{U}_{\lambda}$ is $P$-stable and has dimension coprime to $p$, by [13, Proposition 2.4], $\hat{U}_{\lambda}$ can be extended to $\hat{A}$. This shows (i) and (ii). The statement (iii) just follows from [6, Theorem 1.6]. ∎
As a conesquence, we can get the following decomposition numbers of the block $b$ with nontrivial cyclic hyperfocal subgroup.
@@
With the notation as above, decomposition numbers of the block $b$ are either $p^{l}$ or 0 with $1\leq p^{l}\leq|R/C_{R}(D)|$ when the hyperfocal subgroup $D$ is nontrivial cyclic. In particular, when the hyperfocal subgroup $D$ is nontrivial cyclic and central in $P$, decomposition numbers of the block $b$ are either $1$ or 0.
@@
Since the block algebra $\mbox{$\mathcal{O}$}Gb$ is Morita equivalent to the source algebra $A$, it suffices to show that decomposition numbers of the source algebra $A$ satisfy this property. As before, we can get these numbers by calculating the image $\hat{U}_{A}]$ of $\hat{U}_{A}$ in $R_{\mathcal{K}}(A)$ of any indecomposable projective $A$-module $U_{A}$ such that $\hat{U}_{A}=\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U_{A}$. Since every simple $\bar{\mathbb{A}}$-module can be extended to $\bar{A}$ (see the paragraph 6.16), then every indecomposable projective $A$-module has the form $A\mathfrak{i}$, which is isomorphic to $A\mathop{\otimes}\limits_{\mathbb{A}}\mathbb{A}\mathfrak{i}$ for some primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. Now fix a primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. By Proposition 6.32, we can assume that as $\hat{\mathbb{A}}-$modules, $\hat{\mathbb{A}}\mathfrak{i}\cong\hat{U}_{\lambda}\oplus\hat{U}_{\lambda^{\prime}}$ for some suitable $\lambda,\lambda^{\prime}\in\mathrm{Irr}(E)$ or $\hat{\mathbb{A}}\mathfrak{i}\cong\hat{U}_{\mu}\oplus(\mathop{\bigoplus}\limits_{\chi\in\mathcal{M}}\hat{U}_{\chi})$ for some suitable $\mu\in\mathrm{Irr}(E)$. Fix these notation. Then as $\hat{A}-$modules, $\hat{A}\mathfrak{i}\cong\hat{A}\mathop{\otimes}\limits_{\hat{\mathbb{A}}}\mathbb{A}\mathfrak{i}\cong\mathrm{Ind}_{1}^{R}(\hat{U}_{\lambda})\oplus\mathrm{Ind}_{1}^{R}(\hat{U}_{\lambda^{\prime}})$ or $\hat{A}\mathfrak{i}\cong\hat{A}\mathop{\otimes}\limits_{\hat{\mathbb{A}}}\mathbb{A}\mathfrak{i}\cong\mathrm{Ind}_{1}^{R}(\hat{U}_{\mu})\oplus(\mathop{\bigoplus}\limits_{\chi\in\mathcal{M}}\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi}))$. By Proposition 6.34, as $\hat{A}-$modules, $\mathrm{Ind}_{1}^{R}(\hat{U}_{\nu})\cong\mathop{\bigoplus}\limits_{\vartheta\in\mathrm{Irr}(R)}\hat{U}_{\nu}^{\mathrm{ext}}\vartheta$ for any $\nu\in\mathrm{Irr}(E)$ and $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi})\cong\mathop{\bigoplus}\limits_{\mu_{\chi}\in\mathrm{Irr}(R_{\chi})}\mathrm{Ind}_{R_{\chi}}^{R}(\hat{U}_{\chi}^{\mathrm{ext}}\mu_{\chi})$ for any $\chi\in\mathcal{M}$. When two elements $\chi$ and $\chi^{\prime}$ of $\mathcal{M}$ belong to the same $R$-orbit, the two $\hat{A}-$modules $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi})$ and $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi^{\prime}})$ are isomorphic to each other. Since $C_{R}(D)$ acts trivially on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$, it is easy to get the information on decomposition numbers of $A$ as stated in this corollary. ∎

### v2

@@ 6.2 Cyclic group case
Since $\hat{\mathbb{A}}$ is semi-simple, we have $\hat{U}\cong\hat{U}_{1}\oplus\hat{U}_{2}$ as $\hat{\mathbb{A}}-$modules. Then by [15, Theorem 4.16.4], there is an $\mathcal{O}$-pure submodule $U_{1}$ of $U$ such that $\hat{U}_{1}$ is isomorphic to $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ as $\hat{\mathbb{A}}-$modules. By [15, Proposition 4.2.6], $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ can be viewed as a submodule of $k\mathop{\otimes}\limits_{\mathcal{O}}U$ which is a projective indecomposable $\bar{\mathbb{A}}$-module. Since $\mathbb{A}$ is a symmetric $\mathcal{O}$-algebra, the socle of $k\mathop{\otimes}\limits_{\mathcal{O}}U$ is simple and $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ can not be projective due to the indecomposability of $k\mathop{\otimes}\limits_{\mathcal{O}}U$. At the same time, this implies that the socle of $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ is also simple. In particular, $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ is indecomposable. By Lemma 6.32, $\hat{U}_{1}]$ is not contained in $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$. ∎
@@
Due to the structure of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ described in Proposition 6.26 (iii), by Lemmas 6.32 and 6.33, the proof is similar to the one of [16, Theorem 11.10.5]. ∎
6.35.
Now we will give an explicit description of $\mathrm{Irr}_{\mathcal{K}}(A)$ by the Clifford theoretic relationship between the representation theory of the source algebra $A$ and the hyperfocal subalgebra $\mathbb{A}$. This description can be obtained once it can be proved that the isometry $\Phi_{\mathcal{K}}$ in the commutative diagram (6.25.5) commutes with the conjugation actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}(L)$. This fact seems to be established by directly applied [18, Theorem 2.6] to our case since the isometry $\Phi_{\mathcal{K}}$ is induced by the $P$-invariant Rickard complex $\mathbb{M}_{\cdot}$. However, the validity of Theorem 2.6 in [18] depends on an assumption that the symmetrizing forms of $A$ and $\mbox{$\mathcal{O}$}(P\rtimes E)$ are $R$-invariant symmetrizing forms for $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, which means that there is an $\mathbb{A}$-$\mathbb{A}$-bimodule ($\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodule, resp.) isomorphism between $\mathbb{A}$ ($\mbox{$\mathcal{O}$}L$, resp.) and $\mathbb{A}^{*}$ ($(\mbox{$\mathcal{O}$}L)^{*}$, resp.) commuting with the conjugation action of $R$ (see [17, 5.1.A]). This assumption obviously holds for $\mbox{$\mathcal{O}$}L$ and $\mbox{$\mathcal{O}$}(P\rtimes E)$. For $\mathbb{A}$ and $A$, the proof of Theorem 1.1 in [10] implicitly demonstrates the validity of this assumption (see [10, Section 4]). In conclusion, the isometry $\Phi_{\mathcal{K}}$ in the commutative diagram (6.25.5) commutes with the conjugation actions of $P$.
6.36.
@@
By the arguments in the paragraph 6.35, the isometry $\Phi_{\mathcal{K}}$ in the commutative diagram (6.25.5) preserves the actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}(L)$. Then the statements on stabilizers of simple $\hat{\mathbb{A}}-$modules are clear. For any $\chi\in\mathcal{M}_{R}$, since $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is coprime to $p$, the same argument in the proof of Proposition 6.17 shows that $\hat{U}_{\chi}$ can be extended to $\hat{\mathbb{A}}_{R_{\chi}}$. For any $\lambda\in\mathrm{Irr}(E)$, since $\hat{U}_{\lambda}$ is $P$-stable and has dimension coprime to $p$, by [12, Proposition 2.4], $\hat{U}_{\lambda}$ can be extended to $\hat{A}$. This shows (i) and (ii). The statement (iii) just follows from [6, Theorem 1.6] or [12, Proposition 2.3]. ∎
As a conesquence, we can get the following decomposition numbers of the block $b$ with a nontrivial cyclic hyperfocal subgroup.
@@
With the notation as above, all decomposition numbers of the block $b$ are either $p^{l}$ or 0 with $1\leq p^{l}\leq|R/C_{R}(D)|$ when the hyperfocal subgroup $D$ is nontrivial cyclic. In particular, when the hyperfocal subgroup $D$ is nontrivial cyclic and central in $P$, all decomposition numbers of the block $b$ are either $1$ or 0.
@@
Since the block algebra $\mbox{$\mathcal{O}$}Gb$ is Morita equivalent to the source algebra $A$, it suffices to show that all decomposition numbers of the source algebra $A$ satisfy this property. As before, we can get these numbers by calculating the image $\hat{U}_{A}]$ of $\hat{U}_{A}$ in $R_{\mathcal{K}}(A)$ of any projective indecomposable $A$-module $U_{A}$ with $\hat{U}_{A}=\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U_{A}$. Since every simple $\bar{\mathbb{A}}$-module can be extended to $\bar{A}$ (see the paragraph 6.18), then every projective indecomposable $A$-module has the form $A\mathfrak{i}$, which is isomorphic to $A\mathop{\otimes}\limits_{\mathbb{A}}\mathbb{A}\mathfrak{i}$ for some primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. Now fix a primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. By Proposition 6.34, we can assume that as $\hat{\mathbb{A}}-$modules, $\hat{\mathbb{A}}\mathfrak{i}\cong\hat{U}_{\lambda}\oplus\hat{U}_{\lambda^{\prime}}$ for some suitable distinct $\lambda,\lambda^{\prime}\in\mathrm{Irr}(E)$ or $\hat{\mathbb{A}}\mathfrak{i}\cong\hat{U}_{\mu}\oplus(\mathop{\bigoplus}\limits_{\chi\in\mathcal{M}}\hat{U}_{\chi})$ for some suitable $\mu\in\mathrm{Irr}(E)$. Fix these notation. Then as $\hat{A}-$modules, $\hat{A}\mathfrak{i}\cong\hat{A}\mathop{\otimes}\limits_{\hat{\mathbb{A}}}\mathbb{A}\mathfrak{i}\cong\mathrm{Ind}_{1}^{R}(\hat{U}_{\lambda})\oplus\mathrm{Ind}_{1}^{R}(\hat{U}_{\lambda^{\prime}})$ or $\hat{A}\mathfrak{i}\cong\hat{A}\mathop{\otimes}\limits_{\hat{\mathbb{A}}}\mathbb{A}\mathfrak{i}\cong\mathrm{Ind}_{1}^{R}(\hat{U}_{\mu})\oplus(\mathop{\bigoplus}\limits_{\chi\in\mathcal{M}}\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi}))$. By Proposition 6.37, as $\hat{A}-$modules, $\mathrm{Ind}_{1}^{R}(\hat{U}_{\nu})\cong\mathop{\bigoplus}\limits_{\vartheta\in\mathrm{Irr}(R)}\hat{U}_{\nu}^{\mathrm{ext}}\vartheta$ for any $\nu\in\mathrm{Irr}(E)$ and $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi})\cong\mathop{\bigoplus}\limits_{\mu_{\chi}\in\mathrm{Irr}(R_{\chi})}\mathrm{Ind}_{R_{\chi}}^{R}(\hat{U}_{\chi}^{\mathrm{ext}}\mu_{\chi})$ for any $\chi\in\mathcal{M}$. When two elements $\chi$ and $\chi^{\prime}$ of $\mathcal{M}$ belong to the same $R$-orbit, the two $\hat{A}-$modules $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi})$ and $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi^{\prime}})$ are isomorphic to each other. Since $C_{R}(D)$ acts trivially on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$, it is easy to get the information on decomposition numbers of $A$ as stated in this corollary. ∎